# Exploring the parameter dependence of atomic minima with implicit differentiation

### v1

@@ IV.1 Implementation of the sparse linear operator as a constrained minimization in LAMMPS
Equation (8) defines a linear operator which can be used in iterative solution of the linear equations (5) and thus to evaluate the implicit derivative $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}$. To fully exploit the efficiencies afforded by the Hessian sparsity, in this section, we detail how $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}$ can be implemented in the massively parallel LAMMPS simulation package. This is achieved through $N_{D}$ constrained minimizations, one for each column $[\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}]_{l}$ , $l\in[1,N_{D}]$ of the implicit derivative. We have established, through a wide range of numerical tests using the full dense solution, that the off-diagonal terms $\nabla^{2}_{\tilde{\bf X}\epsilon}{U}$ in (5) can be neglected when determining the homogeneous term $\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}$, meaning the inhomogeneous term $\nabla_{\Theta_{l}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}$ satisfies
$[\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}]_{l}\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}{U}+[{\bf B}]_{l},\quad l\in[1,N_{D}],$ (14)
where ${\bf B}=\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}\otimes\nabla^{2}_{\epsilon\tilde{\bf X}}{U}+\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}\in\mathbb{R}^{N_{D}\times N\times 3}$. For the linear-in-descriptor SNAP potentials used here, the $\nabla^{2}_{\epsilon{{\bm{\Theta}}}}{U}$ term can be directly accessed as derivatives of descriptors [23] using fix sna/atom and related commands in LAMMPS. For the inverse design applications in this paper, which modified only the interaction parameters of a single solute atom without changes to the supercell strain, we have $\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}={\bf 0}$ and thus ${\bf B}=\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}$. With an initial parameter vector ${{\bm{\Theta}}}$, we then define the modified energy function for a parameter index $l\in[1,N_{D}]$
$U(\tilde{\bf X},{\bf C};{{\bm{\Theta}}})+\alpha[{\bf B}]_{l}\cdot[\tilde{\bf X}-\tilde{\bf X}^{*}_{{\bm{\Theta}}}],\quad l\in[1,N_{D}].$ (15)
This modified energy is simple to implement in LAMMPS through the fix addforce function, with similar ease of implementation in any molecular dynamics package. It is straightforward to show that in the limit $\alpha\to 0$ the minimizer $\tilde{\bf X}^{*}_{l,\alpha}$ of (15) gives the $l$th vector $[\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}_{{{\bm{\Theta}}}}]_{l}$ of the implicit derivative $\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}_{{{\bm{\Theta}}}}$ through
@@
Further details regarding hyperparameter scanning for suitable values of $\alpha$ and comparison against the full dense linear solution employing the Hessian matrix is provided in SM.
@@ IV.3 Ab initio QM-ML calculations
The ab initio reference data for Be segregation to screw dislocations in W was calculated using QM-ML hybrid simulations [4], which couple ab initio and machine learning potentials. Initial structures were obtained with matscipy.dislocation module [33]. The Be segregation calculation used the exact same approach reported in [4] for He segregation. Ab initio forces were evaluated using VASP [36, 37] with 10 $\mathbf{k}$-points along the periodic line direction, with a cutoff energy of 500 eV and a minimization force threshold of 0.01 eV/Å. The machine learning force field was a modified SNAP/MILaDy potential from [8], as detailed in [4]. The QM/ML coupling used a buffer radius of 10 Å, resulting in a total of 246 VASP atoms, of which 168 were in the buffer. We refer the reader to [4] for further details.
@@ 1.1 General expression derivation
${\rm d}{\bf F}({\bf X}^{*}_{{{\bm{\Theta}}}+\delta{{\bm{\Theta}}}};{{\bm{\Theta}}}+\delta{{\bm{\Theta}}})=\delta{\bf X}^{*}_{{\bm{\Theta}}}\nabla_{\bf X}{\bf F}({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})+\delta{{\bm{\Theta}}}\cdot\nabla_{{\bm{\Theta}}}{\bf F}({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})+\mathcal{O}(\delta{{\bm{\Theta}}}^{2})={\bf 0}.$ (20)
@@
Expressing atomic coordinates as ${\bf X}=\tilde{\bf X}{\bf C}$ and splitting the position and mixed Hessians on the scale-coordinate and supercell blocks, one gets the equation (5) from the main text.
@@ 1.2 Homogeneous implicit derivative of strain
In this work, we split the full implicit derivative (equation (22)) onto the inhomogeneous and homogeneous contributions. As explained in the main text (Methods A), the inhomogeneous implicit derivative $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}$ can be computed in LAMMPS with constraint minimization. In this section, we detail our finite difference implementation of the homogeneous part of the implicit derivative, $\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}$.
A stationary configuration of a system of volume $V^{*}_{{\bm{\Theta}}}$ corresponds to zero pressure, i.e. $P(V^{*}_{{\bm{\Theta}}})=-\partial{U}(\tilde{\bf X}^{*}_{{\bm{\Theta}}},{\bf C}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})/\partial V=0$. For isotropic supercell variations, ${\bf C}^{*}_{{\bm{\Theta}}}=\left[1+\epsilon^{*}_{{\bm{\Theta}}}\right]{\bf C}_{0}$, this equation can be reformulated in terms of strain:
@@
In analogy with the previous section (equation (20)), we consider a system upon a parameter variation ${{\bm{\Theta}}}+\delta{{\bm{\Theta}}}$:
@@
We then use the definition of the homogeneous implicit derivative $\delta\epsilon^{*}_{{\bm{\Theta}}}=\nabla_{{\bm{\Theta}}}\epsilon_{{\bm{\Theta}}}^{*}\delta{{\bm{\Theta}}}$:
@@ 1.3 Implicit derivative of strain including homogeneous and inhomogeneous terms
$\nabla_{{\bm{\Theta}}}\epsilon_{{\bm{\Theta}}}^{*}=-\frac{\partial{\bf D}(\epsilon^{*}_{{\bm{\Theta}}})/\partial\epsilon+\nabla_{{\bm{\Theta}}}{\bf X}^{*}_{{\bm{\Theta}}}\cdot\partial{\bf F}/\partial\epsilon}{{{\bm{\Theta}}}\cdot\partial^{2}{\bf D}(\epsilon^{*}_{{\bm{\Theta}}})/\partial\epsilon^{2}},$ (32)
where the force derivative is similarly evaluated with the finite difference approach.
@@ 1.4 Taylor expansion of energy
In this section, we derive the expansion of potential energy of a stationary system, ${U}({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})$, due to a perturbation in parameters ${{\bm{\Theta}}}$. For clarity we will omit $({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})$ from ${U}({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})$ in this section. One has to account for contributions arising from the explicit dependence of ${U}$ on parameters and changes in the stationary configurations ${\bf X}^{*}_{{\bm{\Theta}}}$. Here, we first consider the general atomic coordinates ${\bf X}^{*}$ and later split them on scaled coordinates $\tilde{\bf X}^{*}_{{\bm{\Theta}}}$ and supercell ${\bf C}^{*}_{{\bm{\Theta}}}$.
@@
The term proportional to $\nabla_{\bf X}{U}$ vanishes since ${\bf X}^{*}_{{\bm{\Theta}}}$ is a stationary atomic configuration. For the second-order terms, we express $\delta{\bf X}^{*}_{{\bm{\Theta}}}$ using the implicit derivative: $\delta{\bf X}^{*}_{{\bm{\Theta}}}=\delta{{\bm{\Theta}}}\nabla_{{\bm{\Theta}}}{\bf X}^{*}_{{\bm{\Theta}}}$. Referring to equation (22), we further substitute $\nabla_{{\bm{\Theta}}}{\bf X}^{*}_{{\bm{\Theta}}}$ and get: $\delta{\bf X}^{*}_{{\bm{\Theta}}}=-\delta{{\bm{\Theta}}}\nabla^{2}_{{{\bm{\Theta}}}{\bf X}}{U}\big{[}\nabla^{2}_{{\bf X}{\bf X}}{U}\big{]}^{+}$. The final energy expansion up to terms $\mathcal{O}(\delta{{\bm{\Theta}}}^{3})$ reads:
@@
2 MEMORY AND TIME EFFICIENCY OF IMPLICIT DERIVATIVE EVALUATION
@@ 2.1 Automatic differentiation implementation
Here, we use the LJ random fcc alloy (presented in the main text, Results F) as a test system. We compute the implicit derivative with AD using three approaches: 1) Computing the pseudo-inverse of the Hessian matrix with dense linear algebra solution, called dense. 2) Finding ${\bf X}^{*}_{{{\bm{\Theta}}}+\delta{{\bm{\Theta}}}}$ with minimization (gradient descent method was used in this work) and applying AD to the entire pipeline of functions (potential energy and its derivatives) with jaxopt Python library, called jaxopt. 3) Sparse linear operator technique (Results C in the main text) within the AD framework, called sparse.
@@
For this study, we used the best-in-class, NVIDIA A100 80GB graphic card tailored for scientific computations. To track the usage of the GPU memory for each solver and system size, we used the NVIDIA Management Library, nvml through the Python interface provided by the pynvml package. Figure 6 shows the time (panel a) and memory (panel b) required for inhomogeneous implicit derivative evaluation as a function of number of atoms in the system. The inverse approach shows the best time performance, however, it runs out of 80GB GPU memory at a system of 500 atoms. The second fastest, jaxopt method, allows one to achieve system sizes of up to 2000 atoms before saturating the memory. Lastly, sparse technique, is the most memory efficient and reaches the systems of up to 7000 atoms on a single A100 GPU.
@@ 2.2 Implicit derivative implementation using LAMMPS package
Here, we discuss the time and memory efficiency of the inhomogeneous implicit derivative implementation using the LAMMPS software. The test system is a vacancy in bcc tungsten with SNAP potential (Results G in the main text). We used four CPU nodes with two AMD EPYC 7763 CPUs and 512 GB of RAM per node. We monitored RAM usage with the vmstat tool. We test the efficiency of the dense and sparse approaches presented above. Additionally, we explore the efficiency of the constraint energy minimization approach, called energy, that is described in the main text (Methods A). As seen from Figure 7, the energy approach is the most time- and memory-efficient for large systems. Due to the efficient massive parallelization of the LAMMPS software, this method can be effectively scaled and limited only by the amount of available compute resources.
Figure 7: Efficiency of the LAMMPS implicit derivative implementation approaches. a Wall time and b RAM required to compute the inhomogeneous implicit derivative as a function of the number of atoms in a vacancy in a tungsten system with SNAP potential.
@@ 3.1 Accuracy of position change predictions
In this section, we present a comparison of three methods for evaluating the inhomogeneous implicit derivative implemented within the LAMMPS package. The test system is a vacancy in bcc tungsten and SNAP potential parameters are set according to the Results G section of the main text, equation (13), ${{\bm{\Theta}}}(\lambda,m)=\bar{{{\bm{\Theta}}}}+\lambda({{\bm{\Theta}}}_{m}-\bar{{{\bm{\Theta}}}})$, with a strong perturbation of $\lambda=40$. The predicted position changes are computed as
@@
where the inhomogeneous implicit derivative $\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}_{{\bm{\Theta}}}$ is computed with dense, sparse, or energy methods, $\delta{{\bm{\Theta}}}=\lambda({{\bm{\Theta}}}_{m}-\bar{{{\bm{\Theta}}}})$, and positions $\tilde{\bf X}^{*}$ are obtained through the energy minimization of the system with parameters $\bar{{{\bm{\Theta}}}}$. The true position changes are calculated as follows
@@
where $\tilde{\bf X}^{*}({{\bm{\Theta}}})$ are the minimized positions of a system with parameters ${{\bm{\Theta}}}$.
As demonstrated in Fig. 8a, there is a remarkable agreement between the true and predicted positions, with negligible differences among the three computational methods. Given the low computational cost and memory requirements, the energy method stands out as the optimal choice for inhomogeneous implicit derivative evaluation.
Figure 8: Accuracy of the inhomogeneous implicit derivative solvers implemented in LAMMPS. a Predicted vs true position changes for the vacancy in the bcc tungsten system. b Error of the inhomogeneous implicit derivative prediction of position changes.
3.2 Numerical details on LAMMPS constraint minimization
Here, we detail our constraint energy minimization approach (energy) to compute the inhomogeneous implicit derivative implemented in LAMMPS. As explained in the main text, the implicit derivative corresponding to a parameter index $l$ is obtained as $[\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}_{{{\bm{\Theta}}}}]_{l}=\big{(}\tilde{\bf X}^{*}_{l,\alpha}-\tilde{\bf X}^{*}_{{\bm{\Theta}}}\big{)}/\alpha$. We have found that the optimal values of $\alpha$ should be computed for each parameter ${{\bm{\Theta}}}_{l}$ separately as follows
$\alpha(l)=\frac{\alpha_{0}}{\max{\big{(}|\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}|\big{)}}},$ (38)
where $\alpha_{0}$ is a constant.
Figure 8b presents the dependence of the error of the implicit derivative prediction as a function of the $\alpha_{0}$. The error is computed as

### v2

@@ IV.1 Implementation of the sparse linear operator as a constrained minimization in LAMMPS
Equation (8) defines a linear operator which can be used in iterative solution of the linear equations (5) and thus to evaluate the implicit derivative $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}$. To fully exploit the efficiencies afforded by the Hessian sparsity, in this section, we detail how $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}$ can be implemented in the massively parallel LAMMPS simulation package. This is achieved through $N_{D}$ constrained minimizations, one for each row $[\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}]_{l}$, $l\in[1,N_{D}]$ of the implicit derivative. As follows from equations (5), the inhomogeneous term $\nabla_{\Theta_{l}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}$ satisfies
$[\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}]_{l}\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}{U}+[{\bf B}]_{l}={\bf 0},$ (14)
where ${\bf B}=\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}\otimes\nabla^{2}_{\epsilon\tilde{\bf X}}{U}+\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}\in\mathbb{R}^{N_{D}\times N\times 3}$. For the linear-in-descriptor SNAP potentials used here, the $\nabla^{2}_{\epsilon{{\bm{\Theta}}}}{U}$ term can be directly accessed as derivatives of descriptors [29] using fix sna/atom and related commands in LAMMPS. For the inverse design applications in this paper, which modified only the interaction parameters of a single solute atom without changes to the supercell strain, we have $\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}={\bf 0}$ and thus ${\bf B}=\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}$. With an initial parameter vector ${{\bm{\Theta}}}$, we then define the modified energy function for a parameter index $l$
$U(\tilde{\bf X},{\bf C};{{\bm{\Theta}}})+\alpha[{\bf B}]_{l}\cdot[\tilde{\bf X}-\tilde{\bf X}^{*}_{{\bm{\Theta}}}].$ (15)
This modified energy is simple to implement in LAMMPS through the fix addforce function, with similar ease of implementation in any molecular dynamics package. In the limit $\alpha\to 0$, the minimizer $\tilde{\bf X}^{*}_{l,\alpha}$ of (15) gives the $l^{\rm th}$ vector $[\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}_{{{\bm{\Theta}}}}]_{l}$ of the implicit derivative $\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}_{{{\bm{\Theta}}}}$ through
@@
Further details regarding hyperparameter scanning for suitable values of $\alpha$ and comparison against the full dense linear solution employing the Hessian matrix is provided in supplementary material.
@@ IV.3 Ab initio QM-ML calculations
The ab initio reference data for Be segregation to screw dislocations in W was calculated using QM-ML hybrid simulations [40, 4], which couple ab initio and machine learning potentials. Initial structures were obtained with matscipy.dislocation module [43]. The Be segregation calculation used the same approach as for He segregation reported in [4]. Ab initio forces were evaluated using VASP [44, 45] with 10 $\mathbf{k}$-points along the periodic line direction, with a cutoff energy of 500 eV and a minimization force threshold of 0.01 eV/Å. The machine learning force field was a modified SNAP/MILaDy potential from [8]. The QM/ML coupling used a buffer radius of 10 Å, resulting in a total of 246 VASP atoms, of which 168 were in the buffer.
@@ 1.1 General expression derivation
$\mathrm{d}{\bf F}({\bf X}^{*}_{{{\bm{\Theta}}}+\delta{{\bm{\Theta}}}};{{\bm{\Theta}}}+\delta{{\bm{\Theta}}})=\delta{\bf X}^{*}_{{\bm{\Theta}}}\nabla_{\bf X}{\bf F}({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})+\delta{{\bm{\Theta}}}\cdot\nabla_{{\bm{\Theta}}}{\bf F}({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})+\mathcal{O}(\delta{{\bm{\Theta}}}^{2})={\bf 0}.$ (20)
@@
Expressing atomic coordinates as ${\bf X}=\tilde{\bf X}{\bf C}$ and splitting the position and mixed Hessians on the scaled coordinate, off-diagonal, and supercell blocks, $\big{(}\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}{U}\big{)},\big{(}\nabla^{2}_{\tilde{\bf X}\epsilon}{U}\big{)},\big{(}\nabla^{2}_{\epsilon\epsilon}{U}\big{)}$, one gets the equation (5) from the main text.
@@ 1.2 Homogeneous implicit derivative of strain
In this work, we split the full implicit derivative (equation (22)) onto the inhomogeneous and homogeneous contributions. As explained in the main text (methods A), the inhomogeneous implicit derivative $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}$ can be computed in LAMMPS with constraint minimization, or alternatively, with automatic differentiation (AD), section IID from the main text. Here, we detail our finite difference implementation of the homogeneous part of the implicit derivative, $\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}$.
A stationary configuration of a system of volume $V^{*}_{{\bm{\Theta}}}$ corresponds to zero pressure:
$P(V^{*}_{{\bm{\Theta}}})=-\frac{\partial{U}(\tilde{\bf X}^{*}_{{\bm{\Theta}}},{\bf C}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})}{\partial V}\Bigg{|}_{V^{*}_{{\bm{\Theta}}}}=0.$ (23)
For isotropic supercell variations, ${\bf C}^{*}_{{\bm{\Theta}}}=\left[1+\epsilon^{*}_{{\bm{\Theta}}}\right]{\bf C}_{0}$, this equation can be reformulated in terms of strain:
@@
Note that the descriptor vector does not depend on potential parameters ${{\bm{\Theta}}}$. In analogy with the previous section (equation (20)), we consider a system upon a parameter variation ${{\bm{\Theta}}}+\delta{{\bm{\Theta}}}$:
@@
We then use the definition of the homogeneous implicit derivative, $\delta\epsilon^{*}_{{\bm{\Theta}}}=\delta{{\bm{\Theta}}}\nabla_{{\bm{\Theta}}}\epsilon_{{\bm{\Theta}}}^{*}$:
@@ 1.3 Implicit derivative of strain including homogeneous and inhomogeneous terms
$\nabla_{{\bm{\Theta}}}\epsilon_{{\bm{\Theta}}}^{*}=-\frac{\partial{\bf D}(\epsilon)/\partial\epsilon+\nabla_{{\bm{\Theta}}}{\bf X}^{*}_{{\bm{\Theta}}}\,\partial{\bf F}/\partial\epsilon}{{{\bm{\Theta}}}\cdot\partial^{2}{\bf D}(\epsilon)/\partial\epsilon^{2}}\Bigg{|}_{\tilde{\bf X}^{*}_{{\bm{\Theta}}},\epsilon^{*}_{{\bm{\Theta}}}},$ (33)
where the force derivative over the strain is evaluated with the finite difference approach.
@@ 1.4 Taylor expansion of energy
In this section, we derive the expansion of potential energy of a stationary system, ${U}({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})$, under a potential parameter perturbation. For clarity, we will omit $({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})$ from ${U}({\bf X}^{*}_{{\bm{\Theta}}};{{\bm{\Theta}}})$ in this section. For the expansion, one has to account for contributions arising from the explicit dependence of ${U}$ on parameters ${{\bm{\Theta}}}$, but also from the change in stationary configuration ${\bf X}^{*}_{{\bm{\Theta}}}$. Here, we first consider the general atomic coordinates ${\bf X}^{*}_{{\bm{\Theta}}}$ and later split them on scaled coordinates $\tilde{\bf X}^{*}_{{\bm{\Theta}}}$ and supercell ${\bf C}^{*}_{{\bm{\Theta}}}$.
@@
The term proportional to $\nabla_{\bf X}{U}$ vanishes since ${\bf X}^{*}_{{\bm{\Theta}}}$ is a stationary atomic configuration. For the second-order terms, we express $\delta{\bf X}^{*}_{{\bm{\Theta}}}$ using the implicit derivative: $\delta{\bf X}^{*}_{{\bm{\Theta}}}=\delta{{\bm{\Theta}}}\nabla_{{\bm{\Theta}}}{\bf X}^{*}_{{\bm{\Theta}}}$. We further use the implicit derivative expression, equation (22), and get: $\delta{\bf X}^{*}_{{\bm{\Theta}}}=-\delta{{\bm{\Theta}}}\nabla^{2}_{{{\bm{\Theta}}}{\bf X}}{U}\big{[}\nabla^{2}_{{\bf X}{\bf X}}{U}\big{]}^{+}$. The final energy expansion up to terms $\mathcal{O}(\delta{{\bm{\Theta}}}^{3})$ reads:
@@
2 EFFICIENCY OF IMPLICIT DERIVATIVE EVALUATION SCHEMES
@@ 2.1 Automatic differentiation implementation
Here, we use the LJ random fcc alloy (presented in the main text, section IIF) as a test system. We compute the implicit derivative with AD using three approaches: 1) Computing the pseudo-inverse of the Hessian matrix with dense linear algebra solution, called dense. 2) Finding ${\bf X}^{*}_{{{\bm{\Theta}}}+\delta{{\bm{\Theta}}}}$ with minimization (gradient descent method was used in this work) and applying AD to the entire pipeline of functions (potential energy and its derivatives) with jaxopt Python library, called jaxopt. 3) Sparse linear operator technique (section IIC in the main text) within the AD framework, called sparse.
@@
For this study, we used the best-in-class NVIDIA A100 80 GB graphic card tailored for scientific computations. To track the usage of the GPU memory for each solver and system size, we used the NVIDIA Management Library, nvml through the Python interface provided by the pynvml package. Figure 6 shows the time (panel a) and memory (panel b) required for inhomogeneous implicit derivative evaluation as a function of number of atoms in the system. The inverse approach shows the best time performance, however, it runs out of 80 GB GPU memory at a system of 500 atoms. The second fastest, jaxopt method, allows one to achieve system sizes of up to 2000 atoms before saturating the memory. Lastly, sparse technique, is the most memory efficient and reaches the systems of up to 7000 atoms on a single A100 GPU.
@@ 2.2 Implicit derivative implementation using LAMMPS package
Here, we discuss the time and memory efficiency of the inhomogeneous implicit derivative implementation using the LAMMPS software. The test system is a vacancy in bcc tungsten with SNAP potential (section IIG in the main text). We used four CPU nodes with two MD EPYC 7763 CPUs and 512 GB of RAM per node. We monitored RAM usage with the vmstat tool. We test the efficiency of the dense and sparse approaches presented above for AD. Additionally, we explore the efficiency of the constraint energy minimization approach, called energy, that is described in the main text (section IVA). As seen from Figure 7, the energy approach is the most time- and memory-efficient for large systems. Due to the efficient massive parallelization of the LAMMPS software, this method can be effectively scaled, allowing implicit derivative evaluation for systems of arbitrary size.
Figure 7: Efficiency of LAMMPS implicit derivative implementation approaches. a Wall time and b RAM required to compute the inhomogeneous implicit derivative as a function of the number of atoms in a vacancy in a bcc tungsten system with SNAP potential.
@@ 3.1 Accuracy of position change predictions
In this section, we present a comparison of three methods for evaluating the inhomogeneous implicit derivative, dense, sparse, and energy, as discussed above, implemented within the LAMMPS package. The test system is a vacancy in bcc tungsten and SNAP potential parameters are set according to the section IIG of the main text, equation (13), $\delta{{\bm{\Theta}}}=\lambda({{\bm{\Theta}}}_{m}-\bar{{{\bm{\Theta}}}})$, with a strong perturbation of $\lambda=40$ and one selected potential sample $m$. The predicted position changes are computed as
@@
where the inhomogeneous implicit derivative $\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}_{{\bm{\Theta}}}$ is computed with one of the three method and positions $\tilde{\bf X}^{*}$ are obtained through the energy minimization of the system with the reference potential $\bar{{{\bm{\Theta}}}}$. The true position changes are calculated as follows
@@
where $\tilde{\bf X}^{*}({{\bm{\Theta}}})$ are the minimized positions of a system with parameters ${{\bm{\Theta}}}$ using LAMMPS, the minimum image convention is applied to coordinates.
As shown in Fig. 8a, there is a remarkable agreement between the true and predicted positions, with negligible differences among the three computational methods. Given the low computational cost and memory requirements, the energy method stands out as the optimal choice for inhomogeneous implicit derivative evaluation.
Figure 8: Accuracy of the inhomogeneous implicit derivative solvers implemented in LAMMPS. a Predicted vs true position changes for the vacancy in the bcc tungsten system. b Error of the inhomogeneous position changes prediction for the energy method as a function of the $\alpha_{0}$ parameter.
3.2 Numerical details of LAMMPS constraint energy minimization
Here, we detail our constraint energy minimization approach (called energy) to compute the inhomogeneous implicit derivative implemented in LAMMPS. As explained in the main text, the implicit derivative corresponding to a parameter index $l$ is obtained as $[\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}_{{{\bm{\Theta}}}}]_{l}=\big{(}\tilde{\bf X}^{*}_{l,\alpha}-\tilde{\bf X}^{*}_{{\bm{\Theta}}}\big{)}/\alpha$. We have found that the optimal values of $\alpha$ should be computed for each parameter ${{\bm{\Theta}}}_{l}$ separately as follows
$\alpha(l)=\frac{\alpha_{0}}{\max{\big{(}\big{|}\big{[}\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}\big{]}_{l}\big{|}\big{)}}},$ (39)
where $\alpha_{0}$ is a constant and $\big{[}\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}\big{]}_{l}$ is a row $l$ of the mixed Hessian.
Figure 8b presents the dependence of the error of the implicit derivative prediction as a function of the $\alpha_{0}$ parameter. The error is computed as